# PQA: Zero-shot Protein Question Answering for Free-form Scientific Enquiry with Large Language Models

### v1

@@ 3.3 Dataset Analysis
log value
exact cofactor What is a cofactor of this protein? Zn(2+) The cofactor of this protein is Zn(2+). -
is_enzyme F1
@@
True Yes classification
location F1 Where is this protein located? Membrane
@@
Table 2: Questions and example responses for Biochem-Lite (Section A.4 provides further details on motivation and score computation for each metric). Self-Pika is our best performing model. Underlined values indicate the extracted entities for score calculation.
3.4 Benchmarks
Motivation & Design Criteria:
  1. 1.
  2. 2.
Two-Tiered Benchmarking System:
Alongside the benchmarking questions for the PQA task, we also require a robust strategy to assess the correctness of open-ended answers to them. Three possible assessment strategies could be imagined: (1) Exact matching of statements, (2) Relevant entity extraction, and (3) Comparison using stronger LLMs (xLLMs). While exact statement matching is useful in some research domains, it lacks the rigour required for assessing scientific accuracy, as such we selected entity extraction and xLLM comparison. While xLLM comparison, although ideal, is computationally prohibitive, entity extraction pragmatically balances scientific accuracy with computational efficiency. This balance of performance vs efficiency motivated us to design a two-tiered benchmarking system comprised of a light-weight benchmark employing entity extraction for efficiency, and a rigorous xLLM-based benchmark for scientific fidelity.
Metric ID Question Ground truth Self-Pika’s response
@@ 3.4 Benchmarks
Table 3: Questions and example responses for Biochem-ReAct. Self-Pika is our best performing model.
@@
    Biochem-Lite: Pika’s light-weight benchmarking involves a set of pre-defined, scientifically relevant questions with simple answers extracted for each protein using GPT3.5 during Pika-DS creation. These questions are designed to cover a range of biochemical and functional properties (e.g. binding to cofactros or cellular localization) while evaluating multimodal PQA performance at various levels of (1) information extraction from protein embeddings, (2) cross-modal information processing by the LLM, or (3) the LLM’s ability to generate relevant responses (Table 2 and Section A.4). Although the ground truth answers to these questions are single-worded, for evaluation, questions are presented in free-form, and the model provides an open-ended response. This response is then processed for rule-based entity extraction and scoring using adequate metrics.
    • For mw MALE, the mean-absolute log error (MALE) of the predicted molecular weight (MW) is:
      $MALE=\frac{1}{N}\sum_{i=1}^{N}|\log_{10}(\hat{MW}_{i})/(MW_{i})|$ (1)
    • For exact cofactor, the score is computed as:
    • For location F1, the F1 score is defined as:
      $F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (3)
      The predicted class assignments for the calculation of Precision and Recall are based on the exclusive presence of correct labels (membrane, nucleus, or mitochondrion) in the generated response. Responses lacking or containing multiple labels receive the none class label.
    • For is_enzyme F1, the F1 score is used as defined in Equation 3. Responses are classified as True or False based on the exclusive presence of yes or no, respectively. Any deviation is labeled as none.
    • For binary Q mean F1, we have:
      $F1_{\text{binary Q mean}}=\frac{1}{3}\sum_{q=1}^{3}F1_{q}$ (4)
      where $F1_{q}$ is the F1 score for each binary question with class assignments similar to the is_enzyme F1.
@@
    Biochem-ReAct: While light-weight benchmarking questions offer a general overview of model’s performance in various scientific aspects, they do not represent real-world scenarios where the scientific fidelity of open-ended responses is vital. Therefore, we also selected three biochemical questions surrounding reaction & activity of the proteins that can be subjects of scientific enquiry into novel protein sequences (Table 3). The complexity of the responses to these questions necessitates assessment using advanced LLMs such as GPT3.5 and GPT4, therefore imposing a significant computational cost. As a result, Biochem-ReAct is mainly suitable for end-point benchmarking of the best-performing models.
@@
4 Baselines & Methodology
In this section we outline the baselines, architecture and the training strategy for our proposed solutions to the task of scientific PQA. All baseline models and architectures were implemented using PyTorch Lightning, and the complete codebase is accessible on github.com/EMCarrami/Pika.
4.1 Baselines
The prior research in the subject of multimodal text-protein models has primarily been focused on improving protein embeddings with textual data or for text-guided protein sequence generation (Xu et al., 2023; Zhang et al., 2022). The use of PQA in other studies (Guo et al., 2023) has been limited to knowledge retrieval, which do not provide comparable benchmarking opportunities for scientific PQA, mainly due to the high presence of non-scientific content in the PDB-QA dataset (e.g., author names). Therefore, in the absence of relevant state-of-the-art, here we define four baselines specifically for scientific PQA in the context of Biochem-Lite benchmarking questions (Section 3.1).
  • Random Baseline: Selects a random answer for each question from the pool of all possible answers to the same question within the evaluated dataset/split.
  • MLP: An MLP on protein sequence embeddings of the PLM (Section A.5). Applies to classification or value prediction tasks in Biochem-Lite (Table 2). It sets an upper limit for the information content of PLM embeddings, with any improvements potentially indicating generalization due to the knowledge encapsulated in the LLM.
  • LLM Only: Pre-trained LLM’s base response to Biochem-Lite questions. Phi-2 was used.
  • Pika w/o PLM: Mirrors Self-Pika’s architecture and training, substituting the PLM with simple token embeddings of the protein sequences. It defines the base performance for the Biochem-Lite benchmarks (except for “mw MALE”).
@@ 4.1 Baselines
Models Biochem-Lite Biochem-ReAct
@@
pplx Reaction Domains Cofactor
Random 0.35 0.13 0.49 0.34 0.46 - - - -
Baselines MLP 0.07 - 0.86 0.89 - - - - -
Phi-2 only 4.53 0.03 0.34 0.00 0.31 - - - -
Pika w/o PLM 0.02 0.21 0.7 0.47 0.21 2.30 0.01 0.32 0.82
Cross-Pika (Q) 0.02 0.47 0.89 0.77 0.13 1.69 0.36 0.60 0.94
Our Models Self-Pika (Q) 0.04 0.54 0.89 0.76 0.44 1.98 0.58 0.79 0.97
Self-Pika (Q+S) 0.04 0.57 0.88 0.76 0.57 2.23 0.53 0.75 0.96
@@
Table 4: Performance of Pika models on Biochem-Lite and Biochem-ReAct metrics. Bold values indicate the best score. Standard deviation for all scores $<0.05$ except for those indicated by underline. Biochem-Lite results on val set and from three different seeded training. Biochem-ReAct results on test set and a single seed training. “binary mF1” = “binary Q mean F1”, pplx = perplexity, Q = training with QA data only, Q+S = training with QA and Summary data.
4.2 Models
We propose two robust multimodal architectures for PQA tasks, drawing inspiration from successful VLM strategies like Frozen (Tsimpoukelli et al., 2021), Flamingo (Tsimpoukelli et al., 2021) and Prismer (Liu et al., 2023). In both architectures we keep the pre-trained LLM and PLM frozen, only incorporating trainable “adapter” modules that extract relevant latent representations from PLM’s output (protein sequence embeddings) and integrate them into the LLM’s transformer architecture, facilitating seamless transfer of information across modalities.
  • Cross-Pika uses multiple independent learnable adapters, each creating distinct protein latent embeddings for each transformer layer of the LLM. These embeddings are injected into the LLM using a gated cross-attention mechanism before each native self-attention layer (Fig. 3). At the cost of increased complexity, this design allows for a nuanced and layer-specific modulation of the LLM, potentially enabling it to process complex protein-related information more effectively, which could aid the model in answering biochemically intricate PQA queries.
  • Self-Pika uses a single learnable adapter to transform protein sequence embeddings into latent embeddings compatible with language token embeddings. The transformed embeddings, concatenated at the beginning of the LLM’s initial token embeddings, allow the protein latent embeddings to influence the LLM’s response through its internal self-attention mechanisms. This approach effectively creates soft-prompts conditioned on the input protein sequence (Fig. 3). This architecture simplifies the integration process, reducing computational overhead while allowing for the incorporation of essential protein characteristics into the LLM’s response.
Considering the diversity in protein lengths, we opted for the Perceiver architecture (Jaegle et al., 2021) as the learnable adapter for both models. This choice standardizes the transformation of protein embeddings into a consistent number of latent embeddings for seamless cross-modality information transfer. Optionally, predefined placeholder tokens can also be added to each question to signaling the LLM about the presence of the protein sequence. In this work we used ESM2 (Lin et al., 2022) as the pre-trained PLM and Phi-2 (Microsoft, 2023) or GPT2 (Radford et al., 2019) as the pre-trained LLM. Our best performing model is based on the Self-Pika architecture, with ESM2-L (esm2_33t_650M) and Phi-2 pre-trained models.
4.3 Training
We devised a simple training strategy for PQA models, keeping both the PLM and LLM frozen and optimizing for the causal language model loss with AdamW. Hyperparameters were optimised following a greedy search as detaile in section A.6.
All training was performed on a single A100-80GB or H100-80GB Nvidia GPU. Unless otherwise stated, the QA section of the Pika-DS was used for the training of all Pika models. Examples were split to train, validation and test set based on the UniRef50 cluster of their respective protein sequences. As a result, protein sequence from the evaluation sets (validation and test) are never present in training data.
@@ 4.3 Training
Table 5: Perfomance of Self-Pika in the absence of Q/As during training. Bold values indicate best score. Standard deviation for all scores $<0.05$ except for those indicated by underline. S = Summary, Ctrl Q = Control Question, QA = Q/A pairs.
5 Experiments & Results
We start this section by validating our benchmarking strategy, followed by the zero-shot evaluation of best models. In the end, we present the results of our ablation study, highlighting the key factors in scientific PQA model performance.
5.1 Benchmarking Results
Biochem-Lite Results:
Table 4 summarizes the benchmarking results of Pika models in comparison with various baselines. It is evident that both Self- and Cross-Pika architectures outperform all lower-bound baselines, indicating successful cross-modality information transfer to the LLM. Notably, our top models also surpass the upper-bound MLP baseline on key Biochem-Lite benchmarking questions, underscoring the potential for generalization through the LLM’s knowledge. Additionally, we assessed the reliability of ROUGE scores for evaluating PQA model performance. ROUGE scores remained consistent, regardless of whether Self-Pika utilized simple protein token embeddings (w/o PLM) or embedded protein sequences (Figure B.3). This finding confirms the inadequacy of traditional linguistic metrics for the scientific PQA task, thus demonstrating the utility of our Biochem-Lite metrics.
Biochem-ReAct Results:
Our Biochem-ReAct benchmarking, aimed at measuring scientific fidelity, yielded promising results (Table 4). The Self-Pika model demonstrated strong performance across all three Biochem-ReAct questions, exceeding the without PLM baseline by 57%, 31%, and 15% for correctly identifying Reactions, Domains, and Cofactors, respectively, for previously unseen protein sequences. Additionally, comparing Biochem-Lite results against Biochem-ReAct scores for 12 high-performing checkpoints with varied configurations revealed that while traditional linguistic metrics, including perplexity, fall short in predicting the real-world efficacy of PQA models, two Biochem-Lite questions—“exact cofactor recall” and “binary Q mean F1”—emerged as strong indicators of multimodal PQA models’ real-world performance (Figure B.4).
Overall, these results validate our benchmarking strategy and establish Self-Pika as the initial state-of-the-art for the task of scientific PQA.
5.2 Zero-shot PQA:
We define zero-shot PQA as answering novel questions about unseen protein sequences. Given the unbiased nature of Pika-DS and data splits based on UniRef50 IDs, our validation and test sets largely comprise unseen protein sequences. Moreover, the “binary Q” component in Biochem-Lite consists of questions never included in the Pika-DS, suggesting our benchmarking broadly represents zero-shot PQA tasks. This section delves into stricter zero-shot scenarios to better understand PQA model behavior.
PQA Learning without Questions:
To understand the significance of Q/A pairs in the training set, we conducted training under two stringent conditions. The first involved using only summary sentences as labels for proteins without any Q/As. The second condition included summary sentences and a single Control Question, “Is this a real protein?”. During training, we randomly shuffled the tokens within the sequence of half the proteins, setting the expected answer to this question as No (for summary labels all protein sequences remained unchanged). This aims to train the model to understand the task of question answering, with summaries providing the scientific context for learning. Remarkably, introducing the Control Question, even when no other questions were present during training, significantly enhanced performance (Table 5). Additionally, we observed that the performance of the model on the Control Question follows a similar pattern as other metrics (Table 6). These observation suggest that the current bottleneck in PQA performance is likely the LLM and extending this work to larger LLMs could further improve performance and generalization capabilities.
Random
LLM
only
self-Pika
w/o PLM
self-Pika
F1 Score 0.49 0.00 0.43 0.99
Accuracy 0.50 0.00 0.51 0.99
Table 6: Performance on Control Question.

### v2

@@ 3.2.2 Two-Tiered Benchmarking System
exact cofactor $\uparrow$ What is a cofactor of this protein? Zn(2+) The cofactor of this protein is Zn(2+).
is_enzyme F1 $\uparrow$
@@
True Yes
location F1 $\uparrow$ Where is this protein located? Membrane
@@
Table 1: Questions and example responses for Biochem-Lite (See A.6 for our motivation for each metric). Underlined indicates extracted entities for score calculation. $\uparrow$: Higher values better. $\downarrow$: Lower values better.
Metric ID Question Ground Truth Example Response
@@
Table 2: Questions and example responses for Biochem-ReAct. Presented in order of difficulty, with Reaction being the most difficult question for a multimodal LLM to answer based on a protein sequence.
@@
    Biochem-Lite: Pika’s light-weight benchmarking involves a set of pre-defined, scientifically relevant questions with simple answers extracted for each protein using GPT3.5 during Pika-DS creation (Table 1). These questions are designed to cover a range of biochemical and functional properties (e.g. binding to cofactros or cellular localization) while evaluating multimodal PQA performance at various levels (i.e., information extraction from protein embeddings, cross-modal information processing by the LLM, or the LLM’s ability to generate relevant responses). Although the ground truth answers to these questions are single-worded, for evaluation, the questions are presented in free-form, and the model provides open-ended responses, which is processed via rule-based entity extraction and is scored using adequate metrics as detailed in Section A.7.
@@
    Biochem-ReAct: While light-weight benchmarking questions offer a general overview of model’s performance in various scientific aspects, they do not represent real-world scenarios where the scientific fidelity of open-ended responses is vital. Therefore, we also selected three biochemical questions at three distinct difficulty levels surrounding Reaction & Activity of the proteins that can be subjects of scientific enquiry into novel protein sequences (Table 2). The complexity of the responses to these questions necessitates assessment using advanced LLMs such as GPT4. To ensure the high reliability of these assessments we performed an iterative prompt optimization with human feedback following the approach described in Section A.4.
@@
3.3 Pika Baselines
As discussed in Section 2, previous PQA studies have either been limited to knowledge retrieval or have ignored the evolutionary context of protein sequences. Combined with the use of unsuitable metrics such as BLEU, these studies do not provide relevant benchmarking opportunities for scientific PQA. Furthermore, unlike natural multimodal LLMs (e.g., Vision Question Answering) where human-baselines are possible, in PQA it is not possible to determine the upper-bound of performance by comparison to expert evaluations, because humans are unable to extract any information from protein sequences. Therefore, in the absence of relevant baselines or human evaluation for PQA, we introduce two multimodal PQA-LLMs and define various lower- and upper-bound baselines.
@@ 3.3 Pika Baselines
3.3.1 Pika Models
We propose two robust multimodal architectures for PQA tasks, drawing inspiration from successful Vision Language Model (VLM) strategies, replacing the vision-encoder with a protein language model (PLM) (Fig. 3).
  1. 1.
    Cross-Pika is inspired from Flamingo (Tsimpoukelli et al., 2021) and Prismer (Liu et al., 2023). It uses multiple independent learnable adapters, each creating distinct protein latent embeddings for each transformer layer of the LLM. These embeddings are injected into the LLM using a gated cross-attention mechanism before each native self-attention layer. At the cost of increased complexity, this design allows for a nuanced and layer-specific modulation of the LLM, potentially enabling it to process complex protein-related information more effectively, which could aid the model in answering biochemically intricate PQA queries.
  2. 2.
    Self-Pika is based on the architecture proposed in Frozen (Tsimpoukelli et al., 2021) and uses a single learnable adapter to transform protein sequence embeddings into latent embeddings compatible with language token embeddings. The transformed embeddings, concatenated at the beginning of the LLM’s initial token embeddings, allow the protein latent embeddings to influence the LLM’s response through its internal self-attention mechanisms. This approach effectively creates soft-prompts conditioned on the input protein sequence. This architecture simplifies the integration process, reducing computational overhead while allowing for the incorporation of essential protein characteristics into the LLM’s response.
Considering the diversity in protein lengths, we opted for the Perceiver architecture (Jaegle et al., 2021) as the learnable adapter for both models. This choice standardizes the transformation of protein embeddings into a consistent number of latent embeddings for seamless cross-modality information transfer. In both architectures we keep the pre-trained LLM and PLM frozen, only incorporating trainable "adapter" modules that extract relevant latent representations from PLM’s output (protein sequence embeddings) and integrate them into the LLM’s transformer architecture, facilitating seamless transfer of information across modalities. In this work we use ESM2 (Lin et al., 2022) as the pre-trained PLM and Phi-2 (Microsoft, 2023) as the pre-trained LLM (In some experiments GPT2 (Radford et al., 2019) was used, where indicated).
3.3.2 Lower-bound Baselines
  • Random Baseline: Selects a random answer for each question from the pool of relevant answers.
  • LLM Only: Pre-trained Phi-2 model to determine LLM’s base response to Biochem-Lite questions.
  • Pika w/o PLM: Mirrors Self-Pika’s architecture and training, substituting the PLM with simple token embeddings of the protein sequences. This effectively reduces the information to amino acid content only, ignoring the context and order of the sequence.
3.3.3 Upper-bound Baselines
  • MLP: An MLP on protein sequence embeddings of ESM2 (Details in Section A.8). It defines the upper limit for the information content of ESM2 embeddings, with any performance beyond this by a model potentially indicating generalization due to the knowledge encapsulated in the LLM.
  • BLAST: Considering the evolutionary context of protein sequences, for any queried sequence, we identify the closest related protein sequence in the training data using BLAST (Camacho et al., 2009) and return the respective information as the prediction. This is a known strong upper-bound baseline, as homology is a strong predictor of protein’s properties.
4 Experiments & Results
In this section we share our key experimental results. Training details can be found in Sections A.9 and A.10.
Figure 4: Evaluating the effectiveness of Biochem-Lite vs traditional linguistic metrics for scientific accuracy of PQA. Statistical significance is determined through a one-tailed paired t-test across three randomly seeded data subsets and model training (significance guide: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$).
4.1 Validation of Biochem-Lite
First, we assessed the reliability of Biochem-Lite metrics in assessing PQA models performance. To We compared a simple Pika model (Self-Pika with ESM2 + GPT2) against its truncated version, Pika w/o PLM (Section 3.3.2), where ESM2 model was replaced with simple protein token embeddings. Considering that the removal of ESM2 should eliminate all functional information, leaving behind only the amino acid content, we expect that the truncated model must perform significantly worse than the original model. Therefore, metrics that fail to show a significant difference between these two would be unsuitable for assessing PQA models. Our results indicate that Biochem-Lite metrics are significantly better in Self-Pika vs Pika w/o PLM, while all ROUGE metrics fail to highlight any differences (Fig. 4). The behaviour of the mw MALE metric is expected as the size of a protein only depends on its amino acid content. This finding confirms the inadequacy of traditional linguistic metrics for scientific PQA task, demonstrating the utility and importance of our Biochem-Lite metrics.
4.2 Zero-shot PQA:
We define zero-shot PQA as answering questions about unseen protein sequences. Considering the debiased nature of Pika-DS, this can easily be achieved by splitting the data based on UniRef50 ID of each sequence, thereby ensuring that the validation/test sequences have no more than 50% sequence similarity with any of the training sequences. While this ensures the unseen sequence criteria, when considering the evolutionary context of proteins, the 50% similarity threshold is a lenient cutoff. This is because homologous proteins may exhibit as low as 15% sequence similarity (Leander et al., 2022). As a result, on UniRef50-based splits, model performance will be a combination of generalization based on the sequence and model’s ability to identify distant evolutionary relations to training data. While such behaviour may be desired for some PQA applications, to extend the scope, we created a more stringent splitting strategy that focuses the evaluation of model’s generalization primarily on the basis of proteins sequences without the impact of evolutionary relationships. To achieve this we minimized the the evolutionary connection of validation/test sequences to the training data. This was achieved by first grouping sequences into evolutionary groups (EvoGroups) (See Section A.11 for details) followed by splitting the data based on their EvoGrop assignments, ensuring that sequences in validation and test do not have closely related evolutionary counterparts in the training data. Naturally, we expect that the BLAST baseline should show a strong performance on UniRef50-based splits, while its performance should diminish on EvoGroup-based splits.
Baseline or Model Biochem-Lite Biochem-ReAct
@@ 4.2 Zero-shot PQA:
Reaction Domains Cofactor
Random 0.35 (0.01) 0.13 (0.02) 0.49 (0.00) 0.34 (0.02) - - - -
Lower-bound Phi-2 only 4.53 (0.01) 0.03 (0.01) 0.34 (0.00) 0.00 (0.00) 0.31 (0.00) - - -
Pika w/o PLM 0.02 (0.01) 0.21 (0.01) 0.7 (0.01) 0.47 (0.01) 0.21 (0.08) 0.01 0.32 0.82
Upper-bound MLP 0.07 (0.01) - 0.86 (0.01) 0.89 (0.03) - - - -
BLAST - 0.71 (NA) 0.88 (NA) 0.93 (NA) - 0.94 0.89 0.99
Pika Models Cross-Pika 0.02 (0.00) 0.47 (0.15) 0.89 (0.02) 0.77 (0.01) 0.13 (0.02) 0.36 0.60 0.94
Self-Pika 0.04 (0.02) 0.54 (0.04) 0.89 (0.01) 0.76 (0.02) 0.44 (0.04) 0.58 0.79 0.97
@@
Table 3: Performance of Pika models on Uniref50 splits. Bold values indicate the best score, underline indicates second best score. Values in ( ) indicate standard deviation. Biochem-Lite results on val set and from three different seeded training. Biochem-ReAct results on test set and a single seed.
4.2.1 Performance on UniRef50 Splits
Table 3 summarizes the benchmarking results of Pika models on a random UniRef50-based split in comparison with various baselines. It is evident that both Pika architectures outperform all lower-bound baselines, indicating successful cross-modality information transfer to the LLM. Notably, our top model surpasses the upper-bound MLP baseline on key Biochem-Lite benchmarking questions, underscoring the potential for generalization via the LLM’s knowledge. However, as expected, the strong BLAST baseline outperforms all models in most Biochem-Lite and all Biochem-ReAct metrics. Nevertheless, these results indicate that our Pika models based on the modestly sized Phi-2 LLM, are able to comprehend distant evolutionary relations form the context of sequence embeddings.
Baseline or Model Biochem-Lite Biochem-ReAct
exact
cofactor
is_enzyme
F1
location
F1
Reaction Domains Cofactor
BLAST Baseline 0.21 0.72 0.51 0.52 0.49 0.94
Self-Pika Model 0.57 0.85 0.69 0.09 0.54 0.87
Table 4: Performance of Pika model on EvoGroup splits. Bold values indicate the best score. Biochem-Lite results on val set and Biochem-ReAct results on test set both with a single seed.
4.2.2 Performance on EvoGroup Splits
To assess generalization by Pika models, in isolation from the effects of evolutionary relationship of proteins, we performed Self-Pika training on EvoGroup splits and compared the results with BLAST baseline on these splits (Table 4). As expected, our results show a significant reduction in performance of BLAST baseline, specially in more complex metrics such as Domains and Reaction, and to a lower extent in simpler metrics such as Cofactor. This is because, cofactor binding, for instance, relies on small protein motifs that are more conserved across distant evolutionary relations. Crucially, Self-Pika, except on the most complex Reaction metric, retained the majority of its performance on EvoGroup splits, allowing it to surpass the performance of BLAST baseline in most metrics. These observations indicate that multimodal PQA models are capable of inferring functional properties both based on the sequence of proteins as well as using distant evolutionary relations.
4.2.3 Correlation of Biochem Metrics
The Self-Pika model demonstrated strong performance across all three Biochem-ReAct questions on the UniRef split, exceeding the without PLM baseline by 57%, 47%, and 15% for correctly identifying Reactions, Domains, and Cofactors, respectively, for previously unseen protein sequences. This prompted us to study the correlation of Biochem-Lite metrics to Biochem-ReAct metrics. Comparing Biochem-Lite results against Biochem-ReAct scores for 12 high-performing checkpoints with varied configurations revealed that while traditional linguistic metrics, including perplexity, fall short in predicting the real-world efficacy of PQA models, two Biochem-Lite questions, "exact cofactor recall" and "binary Q mean F1", emerged as strong indicators of multimodal PQA models’ real-world performance based on Biochem-ReAct metrics (Fig. B.3).
4.2.4 PQA Learning without Questions
To understand the significance of QA pairs in the training set, we conducted training under two stringent conditions. The first involved using only summary sentences as labels for proteins without any QAs. The second condition included summary sentences and a single Control Question, "Is this a real protein?". During training, we randomly shuffled the tokens within the sequence of half the proteins, setting the expected answer to this question as No (for summary labels all protein sequences remained unchanged). This aims to train the model to understand the task of question answering, with summaries providing the scientific context for learning. Remarkably, introducing the Control Question, even when no other questions were present during training, significantly enhanced performance (Table 5). Additionally, we observed that the performance of the model on the Control Question follows a similar pattern as other metrics (Table B.3). These observation suggest that the current bottleneck in PQA performance is likely the LLM and extending this work to larger LLMs could further improve performance and generalization capabilities.
@@ 4.2.4 PQA Learning without Questions
Table 5: Perfomance of Self-Pika in the absence of QAs during training. Bold values indicate best score. Standard deviation for all scores $<0.05$ except for those indicated by underline. S = Summary, Ctrl Q = Control Question, QA = QA pairs.
4.2.5 PQA Learning with Novel Proteins
4.2.6 Ablation Studies
To identify key contributing factors, we compared model performance across various dimensions. Most notably, reducing the size of the LLM from Phi-2, which has 2.8 billion parameters, to GPT-2 Medium, with 355 million parameters, resulted in a 20% decrease in the key "exact cofactor" metric (Table 6). However, the size of the ESM2 model did not seem to have a significant impact, with performance remaining largely similar between ESM2-S, ESM2-M, and ESM2-L (8M, 65M and 350M parameters, respectively). This observation could be attributed to the richness of representations in the Pre-trained PLMs, potentially highlighting a bottleneck in the LLM or the quality and quantity of the data (Table 6). Concordantly, the comparison of results from the zero-shot experiment also confirms that with a lower volume of data, a tailor-made dataset could yield better results. However, as the dataset size increases—a scenario analogous to the image domain—this factor becomes less crucial. These findings suggest that while LLM capacity significantly influences PQA model performance, the sophistication of sequence embeddings provided by ESMs reaches a point of diminishing returns, underscoring the importance of focusing on LLM enhancements and data quality for future improvements.